Workshop 5 Problems Student Q Search Sheet E 2 A Z OBOU Home Insert Page Layout Formulas Data Review View , Calibri (Body)
Answer #1
  • Requirements

Requirement A

Event

Cash

Liabilities

Common Stock

PIC in Excess

Revenues

Expenses

Net Income

Mar-01

$126,000

$105,000

$21,000

May-02

$240,000

$180,000

$60,000

$366,000

$0

$285,000

$81,000

$0

$0

$0

--Working

Requirement A

Event

Cash

Liabilities

Common Stock

PIC in Excess

Revenues

Expenses

Net Income

Mar-01

=7000*18

=7000*15

=126000-105000

May-02

=12000*20

=12000*15

=240000-180000

  • Other requirements

b.

Common Stock

$285,000

c.

Paid in Capital in Excess of Par

$81,000

d.

Total Paid in Capital

$366,000

e.

Total Assets

$366,000